Quantities r and s vary inversely. When r is 1200, s is 0.35. What is the value of s when r is 2400? Express your answer as a decimal to the nearest thousandth.
For this one, is it just .7? It seems too easy to be just .7 and I don't really know how to do these problems.

Answer by stanbon(75887)

Quantities r and s vary inversely. When r is 1200, s is 0.35. What is the value of s when r is 2400? Express your answer as a decimal to the nearest thousandth.
-----
Equation:
r = k/s
---
Solve for "k" using "When r is 1200, s is 0.35".
1200 = k/0.35
k = 0.35*1200
k = 420
------
Equation:
r = 420/s
----
What is the value of s when r is 2400?
2400 = 420/s
---
s = 420/2400
s = 7/40 = 0.175
=========================
Cheers,
Stan H.
===============